# Wilsonian Renormalization of Neural Network Gaussian Processes

### v1

@@ 3 Renormalizing Gaussian features
$\int$ $\mathcal{D}\varphi_{>}P[\varphi_{>}|\varphi_{<}]e^{-\frac{1}{2\sigma^{2}}\left(\Phi_{>}^{\top}\Phi_{>}+2\Phi_{<}^{\top}\Phi_{>}\right)}=\frac{1}{(2\pi)^{N_{>}/2}}\int\!\mathcal{D}\varphi_{>}e^{-\sum_{k>\kappa}\frac{\varphi^{2}_{k}}{2}}e^{-I_{0}[\Phi_{>}]-I_{\mathrm{int}}[\Phi_{<},\Phi_{>}]}$ (23)
$=e^{-\mathrm{tr}\ln\frac{1}{2}\left(\mathbb{1}_{N_{<}\times N_{<}}+F^{\top}F/\sigma^{2}\right)}\exp\left\{\frac{1}{2\sigma^{4}}\Phi_{<}^{\top}F\left(\mathbb{1}_{N_{<}\times N_{<}}+F^{\top}F/\sigma^{2}\right)^{-1}F^{\top}\Phi_{<}\right\}\leavevmode\nobreak\ ,$
where $F^{\top}F$ is an $N_{>}\!\times\!N_{>}$ matrix with internal (replica) indices contracted, defined via
$I_{0}[\Phi_{>}]+I_{\mathrm{int}}[\Phi_{<},\Phi_{>}]\coloneqq$ $\,\frac{1}{2\sigma^{2}}\sum_{m;kk^{\prime}>\kappa}(f_{mk}-y_{k})(f_{mk^{\prime}}-y_{k}^{\prime})\varphi_{k}\varphi_{k^{\prime}}$ (24)
@@
$\mathrm{tr}$ $\ln\left(\mathbb{1}_{N_{>}\times N_{>}}+F^{\top}F/\sigma^{2}\right)^{1/2}=\mathrm{tr}\left(\frac{FF^{T}}{2\sigma^{2}}-\frac{(FF^{\top})^{2}}{4\sigma^{2}}+\ldots\right)$ (25)
$=\mathrm{tr}\ln\left(\mathbb{1}_{M\times M}+FF^{\top}/\sigma^{2}\right)^{1/2}=\mathrm{tr}\ln\left(\mathbb{1}_{M\times M}+C/\sigma^{2}\right)^{1/2}\leavevmode\nobreak\ ,$
@@
$S_{\mathrm{int}}$ $=-\eta\int\!\mathcal{D}\Phi_{<}P[\Phi_{<}]e^{-\frac{1}{2}\mathrm{tr}\ln\left(C/\sigma^{2}+\mathbb{1}\right)}\exp\left\{\frac{1}{2\sigma^{2}}\Phi_{<}^{\top}\left[\frac{1}{\sigma^{2}}\left(\sigma^{-2}+C^{-1}\right)^{-1}-\mathbb{1}\right]\Phi_{<}\right\}$ (28)
$=-\eta\,\,e^{-\frac{1}{2}\mathrm{tr}\ln\left(C/\sigma^{2}+\mathbb{1}\right)}\!\int\!\mathrm{d}\mu_{x}\,e^{-\frac{1}{2}\left(\bm{f}_{m<}(x)-\bm{y}_{<}(x)\right)^{\top}\left(\sigma^{2}+C\right)^{-1}\left(\bm{f}_{m<}(x)-\bm{y}_{<}(x)\right)}\leavevmode\nobreak\ ,$
where on the second line we have reversed our measure changes above, [^6] Explicitly, $\int\!\mathcal{D}\Phi_{<}P[\Phi_{<}]=\int\!\prod_{k\leq\kappa}\mathrm{d}\varphi_{k}\,P[\varphi_{k}]=\int\!\prod_{k\leq\kappa}\mathrm{d}\varphi_{k}\int\!\mathrm{d}\mu_{x}\,\delta\left(\varphi_{k}-\phi_{k}(x)\right)=\int\!\mathrm{d}\mu_{x}$. and $\bm{f}_{m<}(x)$ and $\bm{y}_{<}(x)$ are $M$-dimensional vectors with elements $f_{m<}(x)$ and $y_{<}(x)$, respectively, cf. (13) and (14). Notably, in this expression all the higher feature modes $\phi_{k>\kappa}$ have been removed, and all higher GP modes $f_{m>}$ enter only through $C$, cf. (26). We note in passing that replacing $C$ with its self-consistent mean-value leads to the same average predictor as in canatar .
@@
$S_{\mathrm{int}}=-\eta\!\int\!\mathrm{d}\mu_{x}\,e^{-\frac{1}{2}\left(\bm{f}_{m<}(x)-\bm{y}_{<}(x)\right)^{\top}\left(\sigma^{2}+C_{y_{>}=0}\right)^{-1}\left(\bm{f}_{m<}(x)-\bm{y}_{<}(x)\right)}\leavevmode\nobreak\ .$ (29)
@@
where $\mathbb{1}_{nm}=\delta_{nm}$. We then set the cutoff such that
@@
In the language of RG, $\sigma^{2}$ is the bare ridge parameter at which the first infinitesimal shell was defined, and $\sigma_{c}^{2}$ is the renormalized value. Of course, this procedure breaks down when the smallest $\lambda$ which remains become comparable to $\sigma_{c}^{2}$. However, long before this point, we will reach the limit $\lambda=\sigma_{c}^{2}/\eta$, which marks the threshold for learnability (cf. the discussion at the end of Sec. 2). Thus, for a generic target function, a natural point to stop the momentum shell RG flow would be at the first learnable mode.
It is interesting to compare this result with other works which use the feature Gaussianity assumption, such as canatar . There it was found that the dataset average predictor coincides with the Equivalent Kernel approximation with an effective ridge parameter, $\kappa$. This is similar to our result in that their $\kappa$ is the original ridge parameter plus all the eigenfunctions which are unlearnable. There are two important conceptual differences however: in that derivation, there is no notion of RG, nor of a Wilsonian cut-off or RG flows. The same largely holds for their concurrent work atanasov2024scaling . Furthermore, the criteria for absorbing unlearnable modes is not sharp as in our case, but is instead controlled by the ratio of $\kappa$ over the number of datapoints and the eigenfunctions.
4 Renormalization beyond the Gaussian
Here we extend the RG formalism to accommodate non-Gaussian effects associated with the feature modes. Namely, cases where the $\varphi$ random variables (under the data measure) have substantial deviations from normality. We begin with a somewhat artificial but illustrative scenario where the greater features are non-Gaussian yet remain independent from the lesser features. In this case, we show that the previous treatment still holds without any changes. We then proceed with allowing weak non-Gaussianity with potential dependencies between greater and lesser feature modes. In this case, alongside the noise renormalization, we obtain a spatial re-weighting of the mean squared error (MSE) loss or, equivalently, a spatial re-weighting of the ridge parameter.
@@ 4 Renormalization beyond the Gaussian
In accordance with our perturbative RG prescription above, we take $\Phi_{>}$ to contain a single feature mode (the rest are to be integrated out in subsequent steps), which we denote by $\varphi_{q}$, so that $\Phi_{q}$ is the replica space vector whose $m^{\mathrm{th}}$ element is $\Phi_{mq}\coloneqq[\Phi_{q}]_{m}=(f_{mq}-y_{q})\varphi_{q}$, cf. (21). The above integral can then be viewed as the cumulant generating function of $P[\varphi_{q}]\exp\{-\frac{1}{2\sigma^{2}}\Phi_{q}^{\top}\Phi_{q}\!\}$ with $-\Phi_{q}^{\top}\Phi_{<}/\sigma^{2}$ playing the role of the cumulant function argument. Accordingly, the leading order non-Gaussian corrections will enter as $(f_{<}(x)-y_{<}(x))^{4}$ times the fourth cumulant of $\varphi_{q}$. For reasonable (e.g., sub-Gaussian) distributions, [^8] Since our approach is perturbative, we limit our study to distributions without heavy tails, cf. zavatoneveth2021exact . the (2n)th cumulant scales like the variance to the $n^{\mathrm{th}}$ power. The higher cumulants are thus $O\left(\delta c^{2}\right)$ and hence negligible, and we obtain the same equations as in the Gaussian case.
We now turn to the more realistic case in which the lesser and greater feature modes are coupled, and relax the Gaussianity assumption by noting that if the data has a high effective dimension, $d$, then a generic feature – which is a sum of many monomials in $x_{i}$ – can be viewed as a sum of many weakly correlated random variables and hence well-approximated by a Gaussian. This, together with the empirical success of the Gaussian assumption canatar , motivates us to consider corrections to Gaussianity as small $1/d$ corrections (as in tomasini2022failure ) and treat the latter perturbatively. At a technical level, we do so by keeping a leading-order Edgeworth expansion of $P[\varphi_{<},\varphi_{>}]$. We then use this expansion to obtain the conditional distribution $P[\varphi_{>}|\varphi_{<}]$ appearing in (22), and perform the integral perturbatively by keeping only the leading non-Gaussianities. Concretely, we take as our large-$d$ expansion
$P[\varphi_{<},\varphi_{>}]\approx{\cal N}[0,\mathbb{1};\varphi_{<}]{\cal N}[0,\mathbb{1};\varphi_{>}]\left(1+\frac{1}{4!}\sum_{k_{1},k_{2},k_{3},k_{4}}U_{k_{1}k_{2}k_{3}k_{4}}He_{4}(\varphi_{k_{1}},\varphi_{k_{2}},\varphi_{k_{3}},\varphi_{k_{4}})\right)$ (35)
@@
$He_{4}(\varphi_{k_{1}},\varphi_{k_{2}},\varphi_{k_{3}},\varphi_{k_{4}})={\cal N}[0,\mathbb{1};\varphi_{<}]^{-1}{\cal N}[0,\mathbb{1};\varphi_{>}]^{-1}\partial_{\varphi_{k_{1}}}\partial_{\varphi_{k_{2}}}\partial_{\varphi_{k_{3}}}\partial_{\varphi_{k_{4}}}{\cal N}[0,\mathbb{1};\varphi_{<}]{\cal N}[0,\mathbb{1};\varphi_{>}]$ (36)
(see for example Naveh_2021 ), and $U$ is the (fourth) cumulant/Ursell function,
$U_{k_{1}k_{2}k_{3}k_{4}}\coloneqq\langle\varphi_{k_{1}}\varphi_{k_{2}}\varphi_{k_{3}}\varphi_{k_{4}}\rangle_{P[\varphi],\mathrm{connected}}$ (37)
i.e., the connected 4-pt correlation function with respect to $P[\varphi]$. We emphasize that in our perturbative expansion in $1/d$, we shall keep only terms which are leading order in the fourth cumulant.
@@
$P[\varphi_{<},\varphi_{q}]$ $\approx{\cal N}[0,\mathbb{1};\varphi_{<}]\left(1+\frac{1}{4!}U_{k1k_{2}k_{3}k_{4}}He_{4}(\varphi_{k_{1}},\varphi_{k_{2}},\varphi_{k_{3}},\varphi_{k_{4}})\right)$ (38)
×$\times{\cal N}[0,\mathbb{1};\varphi_{q}]\left(1+\frac{1}{3!}U_{k_{1}k_{2}k_{3}q}He_{3}(\varphi_{k_{1}},\varphi_{k_{2}},\varphi_{k_{3}})\varphi_{q}+\frac{1}{2!}U_{k_{1}k_{2}qq}He_{2}(\varphi_{k_{1}},\varphi_{k_{2}})[\varphi_{q}^{2}-1]\right)$
@@
$He_{3}(\varphi_{k_{1}},\varphi_{k_{2}},\varphi_{k_{3}})$ $=\varphi_{k_{1}}\varphi_{k_{2}}\varphi_{k_{3}}-\delta_{k_{1},k_{2}}\varphi_{k_{3}}-\delta_{k_{2},k_{3}}\varphi_{k_{1}}-\delta_{k_{1},k_{3}}\varphi_{k_{2}}\leavevmode\nobreak\ ,$
@@
$He_{1}(\varphi_{q})=\varphi_{q}\leavevmode\nobreak\ ,\qquad\qquad He_{2}(\varphi_{q}^{2})=\varphi_{q}^{2}-1\leavevmode\nobreak\ .$ (40)
@@
$P[\varphi_{q}|\varphi_{<}]\equiv\frac{P[\varphi_{q},\varphi_{<}]}{P[\varphi_{<}]}\approx{\cal N}[0,\mathbb{1};\varphi_{q}]\left[1+A\varphi_{q}+B\left(\varphi_{q}^{2}-1\right)\right]\leavevmode\nobreak\ ,$ (41)
@@
$A\coloneqq\frac{1}{3!}U_{k_{1}k_{2}k_{3}q}He_{3}(\varphi_{k_{1}},\varphi_{k_{2}},\varphi_{k_{3}})\leavevmode\nobreak\ ,\quad\quad B\coloneqq\frac{1}{2!}U_{k_{1}k_{2}qq}He_{2}(\varphi_{k_{1}},\varphi_{k_{2}})$ (42)
@@
$\int\!\mathcal{D}\varphi_{>}P[\varphi_{>}|\varphi_{<}]\,e^{-\frac{2\Phi_{<}^{\top}\Phi_{>}+\Phi_{>}^{\top}\Phi_{>}}{2\sigma^{2}}}\approx\int\!\mathcal{D}\varphi_{>}P[\varphi_{>}|\varphi_{<}]\,e^{-\frac{\Phi_{<}^{\top}\Phi_{>}}{\sigma^{2}}}$ (43)
@@
$\int$ $\!\mathcal{D}\varphi_{q}P[\varphi_{q}|\varphi_{<}]e^{-\frac{\Phi_{<}^{\top}\Phi_{q}}{\sigma^{2}}}=\frac{1}{\sqrt{2\pi}}\int\!\mathcal{D}\varphi_{q}e^{-\frac{\varphi^{2}_{q}}{2}-\frac{\Phi_{<}^{\top}\Phi_{q}}{\sigma^{2}}}\left[1+A\varphi_{q}+B\left(\varphi_{q}^{2}-1\right)\right]$ (45)
$=\frac{1}{\sqrt{2\pi}}\int\!\mathcal{D}\varphi_{q}e^{-\frac{\varphi^{2}_{q}}{2}-\frac{1}{\sigma^{2}}\sum_{m}\Phi_{m<}\Phi_{mq}}\left[1+A\varphi_{q}+B\left(\varphi_{q}^{2}-1\right)\right]$
$\approx\exp\left[\frac{1+2B}{2\sigma^{4}}\left(\Phi_{<}^{\top}\bm{f}_{q}\right)^{2}-\frac{A}{\sigma^{2}}\Phi_{<}^{\top}\bm{f}_{q}\right]$
@@
Our next task is to integrate out the greater GP mode ($f_{mq}$), again to leading order in the non-Gaussianity (consistent with our Edgeworth expansion above). Recalling the replicated partition function from Sec. 3, we have
$Z^{M}\right>_{\eta}=e^{-\eta}\int\!\mathcal{D}\bm{f}\,e^{-S_{0}[\bm{f}]+\eta\int\!\mathcal{D}\varphi P[\varphi]\,\exp\left[-\frac{1}{2\sigma^{2}}\left(\Phi_{<}^{\top}\Phi_{<}+2\Phi_{<}^{\top}\Phi_{>}+\Phi_{>}^{\top}\Phi_{>}\right)\right]}$ (46)
@@
$S_{0}[f]=\frac{1}{2}\sum_{m=1}^{M}\sum_{k=1}^{\infty}\frac{1}{\lambda_{k}}|f_{mk}|^{2}\leavevmode\nobreak\ ,$ (47)
@@
$S_{\mathrm{int}}$ $=-\eta\int\!\mathcal{D}\varphi P[\varphi]\,\exp\left[-\frac{1}{2\sigma^{2}}\left(\Phi_{<}^{\top}\Phi_{<}+2\Phi_{<}^{\top}\Phi_{>}+\Phi_{>}^{\top}\Phi_{>}\right)\right]$ (48)
$\approx-\eta\int\!\mathcal{D}\varphi_{<}\,P[\varphi_{<}]\,\exp\left[-\frac{1}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}+\frac{1+2B}{2\sigma^{4}}\left(\Phi_{<}^{\top}\bm{f}_{q}\right)^{2}-\frac{A}{\sigma^{2}}\,\Phi_{<}^{\top}\bm{f}_{q}\right]\leavevmode\nobreak\ .$

### v2

@@ III Renormalizing Gaussian features
$\int$ $\mathcal{D}\varphi_{>}P[\varphi_{>}|\varphi_{<}]e^{-\frac{1}{2\sigma^{2}}\left(\Phi_{>}^{\top}\Phi_{>}+2\Phi_{<}^{\top}\Phi_{>}\right)}$ (25)
$=\frac{1}{(2\pi)^{N_{>}/2}}\int\!\mathcal{D}\varphi_{>}e^{-\sum_{k>\kappa}\frac{\varphi^{2}_{k}}{2}}e^{-I_{0}[\Phi_{>}]-I_{\mathrm{int}}[\Phi_{<},\Phi_{>}]}$
$=e^{-\frac{1}{2}\mathrm{tr}\ln\left(\mathbb{1}_{N_{<}\times N_{<}}+F^{\top}F/\sigma^{2}\right)}$
×$\quad\times\exp\Bigg{\{}\frac{1}{2\sigma^{4}}\Phi_{<}^{\top}F\Big{(}\mathbb{1}_{N_{<}\times N_{<}}+F^{\top}F/\sigma^{2}\Big{)}^{-1}F^{\top}\Phi_{<}\Bigg{\}}\leavevmode\nobreak\ ,$
where $F^{\top}F$ is an $N_{>}\!\times\!N_{>}$ matrix over greater GP modes with internal (replica) indices contracted, such that
$I_{0}[\Phi_{>}]+$ $I_{\mathrm{int}}[\Phi_{<},\Phi_{>}]$ (26)
$\coloneqq$ $\,\frac{1}{2\sigma^{2}}\sum_{m;kk^{\prime}>\kappa}(f_{mk}-y_{k})(f_{mk^{\prime}}-y_{k^{\prime}})\varphi_{k}\varphi_{k^{\prime}}$
@@
$\frac{1}{2}\mathrm{tr}\ln\left(\mathbb{1}_{N_{>}\times N_{>}}+\right.$ $\left.F^{\top}F/\sigma^{2}\right)$ (27)
$=\frac{1}{2}\mathrm{tr}\left(\frac{FF^{T}}{2\sigma^{2}}-\frac{(FF^{\top})^{2}}{4\sigma^{2}}+\ldots\right)$
$=\frac{1}{2}\mathrm{tr}\ln\left(\mathbb{1}_{M\times M}+FF^{\top}/\sigma^{2}\right)$
$=\frac{1}{2}\mathrm{tr}\ln\left(\mathbb{1}_{M\times M}+C/\sigma^{2}\right)\leavevmode\nobreak\ ,$
@@
$S_{\mathrm{int}}$ $=-\eta\int\!\mathcal{D}\Phi_{<}P[\Phi_{<}]e^{-\frac{1}{2}\mathrm{tr}\ln\left(C/\sigma^{2}+\mathbb{1}\right)}$ (30)
×$\times\exp\left\{\frac{1}{2\sigma^{2}}\Phi_{<}^{\top}\left[\frac{1}{\sigma^{2}}\left(\sigma^{-2}+C^{-1}\right)^{-1}-\mathbb{1}\right]\Phi_{<}\right\}$
$=-\eta\,\,e^{-\frac{1}{2}\mathrm{tr}\ln\left(C/\sigma^{2}+\mathbb{1}\right)}\!$
×$\times\int\!\mathrm{d}\mu_{x}\,e^{-\frac{1}{2}\left(\bm{f}_{m<}(x)-\bm{y}_{<}(x)\right)^{\top}\left(\sigma^{2}+C\right)^{-1}\left(\bm{f}_{m<}(x)-\bm{y}_{<}(x)\right)}\leavevmode\nobreak\ ,$
where on the second line we have reversed our measure changes above, [^6] Explicitly, $\int\!\mathcal{D}\Phi_{<}P[\Phi_{<}]=\int\!\prod_{k\leq\kappa}\mathrm{d}\varphi_{k}\,P[\varphi]=\int\!\prod_{k\leq\kappa}\mathrm{d}\varphi_{k}\int\!\mathrm{d}\mu_{x}\,\delta\left(\varphi_{k}-\phi_{k}(x)\right)=\int\!\mathrm{d}\mu_{x}$. and $\bm{f}_{m<}(x)$ and $\bm{y}_{<}(x)$ are $M$-dimensional vectors with elements $f_{m<}(x)$ and $y_{<}(x)$, respectively, cf. (13) and (14). Notably, in this expression all the higher feature modes $\phi_{k>\kappa}$ have been removed, and all higher GP modes $f_{m>}$ enter only through $C$, cf. (28). We note in passing that replacing $C$ with its self-consistent mean-value leads to the same average predictor as in [5].
@@
$S_{\mathrm{int}}=-\eta\!\int\!\mathrm{d}\mu_{x}\,\exp\left\{-\frac{1}{2}\left(\bm{f}_{m<}(x)-\bm{y}_{<}(x)\right)^{\top}\left(\sigma^{2}+C_{y_{>}=0}\right)^{-1}\left(\bm{f}_{m<}(x)-\bm{y}_{<}(x)\right)\right\}\leavevmode\nobreak\ .$ (31)
@@
where $\mathbb{1}_{nm}=\delta_{nm}$. We then set $\delta k$ such that
@@
In the language of RG, $\sigma^{2}$ is the bare ridge parameter at which the first infinitesimal shell was defined, and $\sigma_{c}^{2}$ is the renormalized value. Of course, this procedure breaks down when the smallest $\lambda$ which remains become comparable to $\sigma_{c}^{2}$. However, long before this point, we will reach the limit $\lambda=\sigma_{c}^{2}/\eta$, which marks the threshold for learnability (cf. the discussion at the end of sec. II). Thus, for a generic target function, a natural point to stop the momentum shell RG flow would be at the first learnable mode.
It is interesting to compare this result with other works which use the feature Gaussianity assumption, such as [5]. There it was found that the dataset average predictor coincides with the Equivalent Kernel approximation with an effective ridge parameter, $\kappa$. This is similar to our result in that their $\kappa$ is the original ridge parameter plus all the eigenfunctions which are unlearnable. There are two important conceptual differences however: in that derivation, there is no notion of RG, nor of a Wilsonian cut-off or RG flows. The same largely holds for their concurrent work [7]. Furthermore, the criteria for absorbing unlearnable modes is not sharp as in our case, but is instead controlled by the ratio of $\kappa$ over the number of datapoints and the eigenfunctions.
III.1 Neural Scaling Laws
[…]
IV Renormalization beyond the Gaussian
Here we extend the RG formalism to accommodate non-Gaussian effects associated with the feature modes, namely, cases where the $\varphi$ random variables (under the data measure) have substantial deviations from normality. We begin with a somewhat artificial but illustrative scenario where the greater features are non-Gaussian yet remain independent from the lesser features. In this case, we show that the previous treatment still holds without any changes. We then proceed with allowing weak non-Gaussianity with potential dependencies between greater and lesser feature modes. In this case, alongside the noise renormalization, we obtain a spatial re-weighting of the mean squared error (MSE) loss or, equivalently, a spatial re-weighting of the ridge parameter.
@@ IV Renormalization beyond the Gaussian
In accordance with our perturbative RG prescription above, we take $\Phi_{>}$ to contain a single feature mode (the rest are to be integrated out in subsequent steps), which we denote by $\varphi_{q}$, so that $\Phi_{q}$ is the replica space vector whose $m^{\mathrm{th}}$ element is $\Phi_{mq}\coloneqq[\Phi_{q}]_{m}=(f_{mq}-y_{q})\varphi_{q}$, cf. (23). The above integral can then be viewed as the cumulant generating function of $P[\varphi_{q}]\exp\{-\frac{1}{2\sigma^{2}}\Phi_{q}^{\top}\Phi_{q}\!\}$ with $-\Phi_{q}^{\top}\Phi_{<}/\sigma^{2}$ playing the role of the cumulant function argument. Accordingly, the leading order non-Gaussian corrections will enter as $(f_{<}(x)-y_{<}(x))^{4}$ times the fourth cumulant of $\varphi_{q}$. For reasonable (e.g., sub-Gaussian) distributions, [^8] Since our approach is perturbative, we limit our study to distributions without heavy tails, cf. [38]. the (2n)th cumulant scales like the variance to the $n^{\mathrm{th}}$ power. The higher cumulants are thus $O\left(\delta c^{2}\right)$ and hence negligible, and we obtain the same equations as in the Gaussian case.
We now turn to the more realistic case in which the lesser and greater feature modes are coupled, and relax the Gaussianity assumption by noting that if the data has a high effective dimension, $d$, then a generic feature – which is a sum of many monomials in $x_{i}$ – can be viewed as a sum of many weakly correlated random variables and hence well-approximated by a Gaussian. This, together with the empirical success of the Gaussian assumption [5], motivates us to consider corrections to Gaussianity as small $1/d$ corrections (as in [39]) and treat the latter perturbatively. At a technical level, we do so by keeping a leading-order Edgeworth expansion of $P[\varphi_{<},\varphi_{>}]$. We then use this expansion to obtain the conditional distribution $P[\varphi_{>}|\varphi_{<}]$ appearing in (24), and perform the integral perturbatively by keeping only the leading non-Gaussianities. Concretely, we take as our large-$d$ expansion
$P[\varphi_{<},\varphi_{>}]\approx{\cal N}[0,\mathbb{1};\varphi_{<}]{\cal N}[0,\mathbb{1};\varphi_{>}]$
×$\times\Bigg{(}1+\frac{1}{4!}\sum_{\begin{subarray}{c}k_{1},k_{2},\\ k_{3},k_{4}\end{subarray}}U_{k_{1}k_{2}k_{3}k_{4}}He_{4}(\varphi_{k_{1}},\varphi_{k_{2}},\varphi_{k_{3}},\varphi_{k_{4}})\Bigg{)}\leavevmode\nobreak\ ,\quad$ (37)
@@
$He_{4}$ $(\varphi_{k_{1}},\varphi_{k_{2}},\varphi_{k_{3}},\varphi_{k_{4}})$ (38)
$={\cal N}[0,\mathbb{1};\varphi_{<}]^{-1}{\cal N}[0,\mathbb{1};\varphi_{>}]^{-1}$
×$\quad\times\partial_{\varphi_{k_{1}}}\partial_{\varphi_{k_{2}}}\partial_{\varphi_{k_{3}}}\partial_{\varphi_{k_{4}}}{\cal N}[0,\mathbb{1};\varphi_{<}]{\cal N}[0,\mathbb{1};\varphi_{>}]\leavevmode\nobreak\ ,\quad$
(see for example [40]), and $U$ is the (fourth) cumulant/Ursell function,
$U_{k_{1}k_{2}k_{3}k_{4}}\coloneqq\langle\varphi_{k_{1}}\varphi_{k_{2}}\varphi_{k_{3}}\varphi_{k_{4}}\rangle_{P[\varphi],\mathrm{connected}}\leavevmode\nobreak\ ,$ (39)
i.e., the connected 4-pt correlation function with respect to $P[\varphi]$. We emphasize that in our perturbative expansion in $1/d$, we shall keep only terms which are leading order in the fourth cumulant. Notably the lack of odd cumulants is consistent with our previous simplifying assumption that all features are odd functions therefore $P[\varphi]=P[-\varphi]$. In general, we think of $\varphi$ as spanning some relevant (i.e., potentially learnable) subspace of function space and do not require this subspace to be closed under the multiplication of features (e.g., $\phi_{k}(x)^{2}$, is an even function and thus outside this subspace). A different set of assumptions is taken in sec. (IV.2), leading to similar results.
@@
$P$ $[\varphi_{<},\varphi_{q}]$ (40)
$\approx{\cal N}[0,\mathbb{1};\varphi_{<}]\left(1+\frac{1}{4!}U_{k1k_{2}k_{3}k_{4}}He_{4}(\varphi_{k_{1}},\varphi_{k_{2}},\varphi_{k_{3}},\varphi_{k_{4}})\right)$
×$\times{\cal N}[0,\mathbb{1};\varphi_{q}]\Bigg{(}1+\frac{1}{3!}U_{k_{1}k_{2}k_{3}q}He_{3}(\varphi_{k_{1}},\varphi_{k_{2}},\varphi_{k_{3}})\varphi_{q}$
$\qquad\qquad\qquad\quad+\frac{1}{2!}U_{k_{1}k_{2}qq}He_{2}(\varphi_{k_{1}},\varphi_{k_{2}})[\varphi_{q}^{2}-1]\Bigg{)}\leavevmode\nobreak\ ,$
@@
$He_{3}(\varphi_{k_{1}},\varphi_{k_{2}},\varphi_{k_{3}})$ $=\varphi_{k_{1}}\varphi_{k_{2}}\varphi_{k_{3}}-\delta_{k_{1},k_{2}}\varphi_{k_{3}}$ (42)
$\quad-\delta_{k_{2},k_{3}}\varphi_{k_{1}}-\delta_{k_{1},k_{3}}\varphi_{k_{2}}\leavevmode\nobreak\ ,$
@@
$He_{1}(\varphi_{q})=\varphi_{q}\leavevmode\nobreak\ ,\qquad He_{2}(\varphi_{q}^{2})=\varphi_{q}^{2}-1\leavevmode\nobreak\ .$ (43)
@@
$P[\varphi_{q}|\varphi_{<}]$ $\equiv\frac{P[\varphi_{q},\varphi_{<}]}{P[\varphi_{<}]}$ (44)
$\approx{\cal N}[0,\mathbb{1};\varphi_{q}]\left[1+A\varphi_{q}+B\left(\varphi_{q}^{2}-1\right)\right],$
@@
$A\leavevmode\nobreak$ $\coloneqq\frac{1}{3!}U_{k_{1}k_{2}k_{3}q}He_{3}(\varphi_{k_{1}},\varphi_{k_{2}},\varphi_{k_{3}}),$
$B\leavevmode\nobreak$ $\coloneqq\frac{1}{2!}U_{k_{1}k_{2}qq}He_{2}(\varphi_{k_{1}},\varphi_{k_{2}}),$ (45)
@@
$\int\!\mathcal{D}\varphi_{>}$ $P[\varphi_{>}|\varphi_{<}]\,e^{-\frac{2\Phi_{<}^{\top}\Phi_{>}+\Phi_{>}^{\top}\Phi_{>}}{2\sigma^{2}}}$ (46)
$\approx\int\!\mathcal{D}\varphi_{>}P[\varphi_{>}|\varphi_{<}]\,e^{-\frac{\Phi_{<}^{\top}\Phi_{>}}{\sigma^{2}}}.$
@@
$\int\!\mathcal{D}\varphi_{>}$ $P[\varphi_{>}|\varphi_{<}]\,e^{-\frac{\Phi_{<}^{\top}\Phi_{>}}{\sigma^{2}}}$ (47)
@@
$\int$ $\!\mathcal{D}\varphi_{q}P[\varphi_{q}|\varphi_{<}]e^{-\frac{\Phi_{<}^{\top}\Phi_{q}}{\sigma^{2}}}$ (48)
$=\frac{1}{\sqrt{2\pi}}\int\!\mathcal{D}\varphi_{q}e^{-\frac{\varphi^{2}_{q}}{2}-\frac{\Phi_{<}^{\top}\Phi_{q}}{\sigma^{2}}}\left[1+A\varphi_{q}+B\left(\varphi_{q}^{2}-1\right)\right]$ (49)
$=\frac{1}{\sqrt{2\pi}}\int\!\mathcal{D}\varphi_{q}e^{-\frac{\varphi^{2}_{q}}{2}-\frac{1}{\sigma^{2}}\sum_{m}\Phi_{m<}\Phi_{mq}}$
×$\qquad\qquad\qquad\qquad\times\left[1+A\varphi_{q}+B\left(\varphi_{q}^{2}-1\right)\right]$ (50)
$\approx\exp\left[\frac{1+2B}{2\sigma^{4}}\left(\Phi_{<}^{\top}\bm{f}_{q}\right)^{2}-\frac{A}{\sigma^{2}}\Phi_{<}^{\top}\bm{f}_{q}\right]\leavevmode\nobreak\ ,$ (51)
@@
Our next task is to integrate out the greater GP mode ($f_{mq}$), again to leading order in the non-Gaussianity (consistent with our Edgeworth expansion above). Recalling the replicated partition function from sec. III, we have
$Z^{M}\right>_{\eta}=e^{-\eta}\int\!\mathcal{D}\bm{f}\,\exp\left\{-S_{0}[\bm{f}]+\eta\int\!\mathcal{D}\varphi P[\varphi]\,\exp\left[-\frac{1}{2\sigma^{2}}\left(\Phi_{<}^{\top}\Phi_{<}+2\Phi_{<}^{\top}\Phi_{>}+\Phi_{>}^{\top}\Phi_{>}\right)\right]\right\},$ (52)
@@
$S_{0}[f]=\frac{1}{2}\sum_{m=1}^{M}\sum_{k=1}^{\infty}\frac{1}{\lambda_{k}}|f_{mk}|^{2}\leavevmode\nobreak\ .$
@@
$S_{\mathrm{int}}=$ (53)
$-\eta\int\!\mathcal{D}\varphi P[\varphi]\,\exp\left[-\frac{1}{2\sigma^{2}}\left(\Phi_{<}^{\top}\Phi_{<}+2\Phi_{<}^{\top}\Phi_{>}+\Phi_{>}^{\top}\Phi_{>}\right)\right]$
$\approx-\eta\int\!\mathcal{D}\varphi_{<}\,P[\varphi_{<}]\,$
×$\quad\times\exp\left[-\frac{1}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}+\frac{1+2B}{2\sigma^{4}}\left(\Phi_{<}^{\top}\bm{f}_{q}\right)^{2}-\frac{A}{\sigma^{2}}\,\Phi_{<}^{\top}\bm{f}_{q}\right]\leavevmode\nobreak\ .$